# Graph discretization of Laplacian on Riemannian manifolds with Bounds on Ricci curvature

### v2

@@ 1. Introduction
In the deterministic setting, a major advance made by Burago, Ivanov, and Kurylev [6] established uniform spectral convergence of the graph Laplacian to the Laplace-Beltrami operator for closed Riemannian $n$-manifolds under uniform sectional curvature bounds, a lower injectivity radius bound, and an upper diameter bound. Aubry [2] obtained a similar result on approximating spectral data using isometrically immersed graphs in manifolds. Burago et al. [7] and Lu [18] also extended their framework to the $\rho$-Laplacian on metric measure spaces, including those with boundaries or certain singularities; further generalizations to vector bundles appear in [8]. Some recent related developments include [15], [9], [17], [11] and, [5].
@@
$|\lambda_{k}(\Gamma)-\lambda_{k}(M)|\leq C_{\mathcal{M}}\left(\frac{\epsilon}{\rho}+\rho\right)\lambda_{k}(M)+\rho C_{\mathcal{M}}\lambda^{\frac{3}{2}}_{k}(M),\quad\forall\ k.$
@@
Consequently, for every fix $k$, $\lambda_{k}(\Gamma)$ converges uniformly to $\lambda_{k}(M)$ for all $M\in\mathcal{M}(n,\lambda,D,i_{0})$ as $\rho\to 0$ and $\frac{\epsilon}{\rho}\to 0.$
@@ 2. Some inequalities and Average dispersion
$\exp_{x}^{*}\text{{vol}}_{g}=J(u,t)t^{n-1}dtdu.$
where $J(u,t)$ is called the Jacobian determinant at $(u,t)$. From the proof of the Bishop-Gromov volume comparison theorem, we know that $J(u,0)=1$ for any $u\in S^{n-1}.$
@@
Consider a normal cartesian coordinate on $B_{r}(x)$, the geodesic ball of radius $r$ centered at $x$ in $(M,g)\in\mathcal{M}$. Let the coefficients of the Riemannian metric be $g_{ij}$ with respect to this coordinate. Then the volume element at $p\in B_{x}(r)$ is given by $\tilde{J}_{g}(p)dx_{1}\wedge dx_{2}\wedge..\wedge dx_{n}$ where $\tilde{J}_{g}(p)=\sqrt{det(g_{ij}(p))}.$ Since the metric is smooth, $\tilde{J}_{g}$ is a smooth function on $B_{r}(x)$ and $\tilde{J}_{g}(x)=1.$ For any $u\in S^{n-1}$, let $\tilde{J}^{\prime}_{g}(p)$ denote its partial derivative in the direction $u.$ $\tilde{J}^{\prime}$ is a smooth function on $S^{n-1}\times B_{r}(x)$ and its expression in the given coordinate contains $u,g_{ij}$ and $g_{ij}^{\prime}$. From Anderson’s precompactness theorem in [1], we know that $\mathcal{M}$ is precompact in the $C^{1,\alpha}$ topology. Hence, $u,g_{ij}$ and $g_{ij}^{\prime}$ are bounded for any $M\in\mathcal{M}$. We obtain a uniform upper bound $c_{\mathcal{M}(n,\lambda,D,i_{0})}$ on the derivative on $B_{x}(r)$ for $r\leq\frac{i_{0}}{2}$.
$|\tilde{J}_{g}^{\prime}(p)|\leq c_{\mathcal{M}},\quad\forall\ p\in B_{\frac{i_{0}}{2}}(x),\quad\forall\ M\in\mathcal{M}.$ (2.1)
Using mean value theorem and (2.1), we have
$|\tilde{J}_{g}(p)-1|=|\tilde{J}_{g}(p)-\tilde{J}_{g}(x)|\leq c_{\mathcal{M}}|x-p|$
where |xp||x-p| denotes the distance between $x$ and $p.$ If $x=(u,t)$ in normal polar coordinate then
$|\tilde{J}_{g}(p)-1|=|J(u,t)-1|\leq c_{\mathcal{M}}t.$
@@
Let $\textit{M}\in\mathcal{M}(n,\lambda,D,i_{0})$ and $B_{r}(x)$ be a ball in M centered at $x$ of radius $r$. Let $f:\textit{M}\to\mathbb{R}$ be a smooth function. The average dispersion $E_{r}(f):L^{2}(\textit{M})\to\mathbb{R}$ is defined as
@@
We recall that $H^{1}(M)$ is the space of all functions in $L^{2}(M)$ such that their partial derivatives are also in $L^{2}(M)$. The following lemma gives an upper bound of $E_{r}(f)$.
@@ 3. An upper bound for $\lambda_{k}(\Gamma)$
If $0<\epsilon<r<\rho<\frac{i_{0}}{2}$ and ${\text{vol}}(B_{r}(x))\geq c_{n}r^{n}$ then
@@ 4. Smoothing Operator
We observe that $\Lambda_{r}f=f$, for all constant functions $f.$ In fact, $\Lambda_{r}f$ is bounded when $r$ is less than the injectivity radius as we see in the following lemma. The next lemma gives a bound on the $L^{2}$ norm of the gradient of $\theta$.
@@
Let $0<r<\frac{i_{0}}{2}\text{ and }M\in\mathcal{M}(n,\lambda,D,i_{0})$. For all $f\in L^{2}(M),$ we have
@@
Let $0<r<\frac{i_{0}}{2}\text{ and }M\in\mathcal{M}(n,\lambda,D,i_{0})$. For all $f\in L^{2}(M)$,
@@ 5. A lower bound for $\lambda_{k}(\Gamma)$
Using the above lemma and the results obtained on the interpolation map, we are now ready to compute, under certain constraints, a lower bound for the eigenvalue of the graph Laplacian operator using the eigenvalue of the Laplace-Beltrami operator for all $M\in\mathcal{M}(n,\lambda,D,i_{0})$. Without loss of generality, we consider $\lambda_{\mathcal{M},k}>3(1+4c_{\mathcal{M}}).$
@@ 6. Approximation of eigenfunctions
    Let $u^{\prime}=u-\mathbb{P}_{\lambda+a}u.$ Since $Q(u^{\prime})\geq(\lambda+a)\left\lVert u^{\prime}\right\rVert^{2},$,
@@
The proof of the above theorem is analogous to that of Theorem 4[6] in [6], utilizing the results of Lemma 6.4 and Lemma 6.2. ∎

### v3

# arXiv:2501.18323v3[math.SP] 02 Mar 2026

@@ 1. Introduction
In the deterministic setting, a major advance made by Burago, Ivanov, and Kurylev [8] established uniform spectral convergence of the graph Laplacian to the Laplace-Beltrami operator for closed Riemannian $n$-manifolds under uniform sectional curvature bounds, a lower injectivity radius bound, and an upper diameter bound. Aubry [2] obtained a similar result on approximating spectral data using isometrically immersed graphs in manifolds. Burago et al. [9] and Lu [20] also extended their framework to the $\rho$-Laplacian on metric measure spaces, including those with boundaries or certain singularities; further generalizations to vector bundles appear in [7]. Some recent related developments include [16], [10], [19], [12] and, [6]. Applications of the spectral convergence of graph Laplacians also arise in manifold learning and numerical computation, see for instance [[4],[18],[24],[25]].
@@
$|\lambda_{k}(\Gamma)-\lambda_{k}(M)|\leq C_{\mathcal{M}}\left(\frac{\epsilon}{\rho}+\rho\right)\lambda_{k}(M)+\rho C_{\mathcal{M}}\lambda^{\frac{3}{2}}_{k}(M),\quad\forall\ k\leq|\Gamma|.$
@@
Consequently, for every fixed $k$, $\lambda_{k}(\Gamma)$ converges uniformly to $\lambda_{k}(M)$ for all $M\in\mathcal{M}(n,\lambda,D,i_{0})$ as $\rho\to 0$ and $\frac{\epsilon}{\rho}\to 0.$
We now state the approximation of eigenfunctions of the Laplace-Beltrami operator by using the discretization scheme introduced above. We work with the discretization map $P$ (Definition 3) that connects functions on the manifold with functions on the discrete space and the interpolation map $I$ (Definition 7) which reconstructs a function on the manifold from discrete data.
Theorem 1.3.
$\lambda_{k-1}<\lambda_{k}=\lambda=\lambda_{k+m-1}<\lambda_{k+m}.$
and
@@ 2. Some inequalities and Average dispersion
$\exp_{x}^{*}\operatorname{vol}_{g}=J(u,t)t^{n-1}dtdu.$
where $J(u,t)$ is called the Jacobian determinant at $(u,t)$. In the sequel, we require suitable bounds $\lvert J(u,t)\rvert$ under the assumed geometric conditions. We first establish an estimate for the Laplacian of the distance function under an upper Ricci curvature bound.
Lemma 2.1 (Laplacian comparison of the distance function).
Let $(M,g)$ be a Riemannian n-manifold satisfying $\operatorname{Ric}\leq\lambda$, and $\operatorname{inj}(M)\geq i_{0}>0$. Fix $p\in M$ and let $r(x)=d(p,x)$ be the distance function from $p$. Then at $p\in B_{i_{0}/2}(x)\setminus\{x\}$,
$\Delta r(p)\geq\frac{n-1}{r(p)}-\lambda\,r(p).$
Proof.
Let $\gamma:[0,r(p)]\to M$ be the unique unit-speed minimizing geodesic from $x$ to $p$. One can write in polar coordinates
$g=dr^{2}+g_{r}.$
Let $H_{r}$ denote the Hessian of the distance function $r$. Then the fundamental equation (see [22] Chapter 2, section 5.2) can be written as
$\partial_{r}g=\partial_{r}g_{r}=2H_{r}$
and
$\partial_{r}H_{r}-H_{r}^{2}=-\operatorname{R}(.,\partial_{r},\partial_{r},.).$
By a taking trace and using the curvature assumption $\operatorname{Ric}\leq\lambda$, we have
$\partial_{r}\Delta r\geq\operatorname{tr}(H_{r}^{2})-\lambda\geq\frac{(\Delta r)^{2}}{n-1}-\lambda.$ (2.1)
The last inequality follows from the Cauchy-Schwarz inequality. Therefore,
$\partial_{r}\left(\Delta_{r}-\frac{n-1}{r}\right)$ $=\partial_{r}\Delta r+\frac{n-1}{r^{2}}-2\frac{\Delta r}{r}$
$\geq\frac{1}{n-1}(\Delta r)^{2}-\lambda+\frac{n-1}{r^{2}}-2\frac{\Delta r}{r}$ (2.2)
$=\frac{1}{n-1}\left(\Delta r-\frac{n-1}{r}\right)^{2}-\lambda$
$\geq-\lambda.$ (2.3)
Let $f(r)=\Delta r-\frac{n-1}{r}.$ Then,
$f(r)-f(\epsilon)$ $=\int_{\epsilon}^{r}f^{\prime}(s)ds$
$\geq-\int_{\epsilon}^{r}\lambda ds$
$=-\lambda(r-\epsilon).$ (2.4)
From [22] Chapter 5, section 7
$\lim_{r\to 0}\left(H_{r}-\frac{1}{r}g_{r}\right)=0.$
Taking the trace we have,
$\lim_{\epsilon\rightarrow 0}\left(\Delta r(\gamma(\epsilon))-\frac{n-1}{\epsilon}\right)=0.$
Hence,
$f(r)\geq-\lambda r.$
Which implies,
$\Delta r(p)\geq\frac{n-1}{r(p)}-\lambda r(p).$
@@
Consider a polar coordinate decomposition on $B_{r}(x)$, the geodesic ball of radius $r\leq\frac{i_{0}}{2}$ centered at $x$ in $(M,g)\in\mathcal{M}.$ Then, from [22], Chapter 9, section 1 we have
$d\operatorname{vol}=\tilde{J}(u,t)du\wedge d\operatorname{vol}_{n-1}$
and
$\partial_{r}\tilde{J}=\tilde{J}\Delta r.$ (2.5)
where $\tilde{J}(u,t)=J(u,t)t^{n-1}.$ By differentiating (2.5), we obtain
$\Delta r=\frac{n-1}{r}+\frac{J^{\prime}(u,r)}{J(u,r)}.$ (2.6)
From Lemma 2.1 and the Laplace comparison theorem (see [22], Chapter 9, section 1.2) for $\operatorname{Ric}\geq-\lambda$, we have
$\frac{n-1}{t}-\lambda t\leq\Delta r(\gamma(t))\leq(n-1)\sqrt{\lambda}\coth(\sqrt{\lambda}t).$
Using (2.6),
$\lambda tJ(u,t)\leq J^{\prime}(u,t)\leq(n-1)\sqrt{\lambda}\coth(\sqrt{\lambda}t)J(u,t)-\frac{n-1}{t}J(u,t).$
From the Taylor series expansion of the $\coth$ function, we have $\lim_{t\rightarrow 0}\left(\sqrt{\lambda}\coth(\sqrt{\lambda}t)-\frac{1}{t}\right)=0$ and $J(u,t)\leq\left(\frac{\sinh(\sqrt{\lambda}t}{\sqrt{\lambda}t}\right)^{n-1}.$ Using $|J(u,t)|\leq C(n,\lambda,i_{0})$, we obtain a uniform upper bound $c_{\mathcal{M}(n,\lambda,D,i_{0})}$ on the derivative of $J(u,t)$ on $B_{x}(r)$ for $r\leq\frac{i_{0}}{2}.$
$|J^{\prime}(u,t)|\leq c_{\mathcal{M}},\quad\forall p=\exp_{x}(tu)\in B_{\frac{i_{0}}{2}}(x),\quad\forall M\in\mathcal{M}.$ (2.7)
From the proof of the Bishop-Gromov volume comparison theorem, we know that $J(u,0)=1$ for any $u\in S^{n-1}.$ Using mean value theorem and (2.7) we have
$|J(u,t)-1|=|J(u,t)-J(u,0)|\leq c_{\mathcal{M}}t.$
@@
Let $\textit{M}\in\mathcal{M}(n,\lambda,D,i_{0})$ and $B_{r}(x)$ be a ball in M centered at $x$ of radius $r$. Let $f:\textit{M}\to\mathbb{R}$ be a smooth function. The average dispersion $E_{r}(f)$ is defined as
@@
We recall that $H^{1}(M)$ is the space of all functions in $L^{2}(M)$ such that their first-order partial derivatives are also in $L^{2}(M)$. The following lemma gives an upper bound of $E_{r}(f)$.
@@ 3. An upper bound for $\lambda_{k}(\Gamma)$
If $0<\epsilon<\frac{\rho}{n}<\frac{i_{0}}{2n}$ and ${\text{vol}}(B_{r}(x))\geq c_{n}r^{n}$ for $\epsilon<r<\rho$ then
@@ 4. Smoothing Operator
The next lemma gives a bound on the $L^{2}$ norm of the gradient of $\theta$.
@@
We observe that $\Lambda_{r}f=f$, for all constant functions $f.$ In fact, $\Lambda_{r}f$ is bounded when $r$ is less than the injectivity radius as we see in the following lemma.
@@
Let $0<r<\frac{i_{0}}{2}\text{ and }M\in\mathcal{M}(n,\lambda,D,i_{0})$. For all $f\in L^{2}(M)$ and $r<\frac{1}{c_{\mathcal{M}}}$, we have
@@
Let $0<r<\frac{i_{0}}{2}\text{ and }M\in\mathcal{M}(n,\lambda,D,i_{0})$. For all $f\in L^{2}(M)$ and $r<\frac{1}{c_{\mathcal{M}}}$,
@@ 5. A lower bound for $\lambda_{k}(\Gamma)$
Using the above lemma and the results obtained on the interpolation map, we are now ready to compute, under certain constraints, a lower bound for the eigenvalue of the graph Laplacian operator using the eigenvalue of the Laplace-Beltrami operator for all $M\in\mathcal{M}(n,\lambda,D,i_{0})$. Since $\lambda_{\mathcal{M},k}$ is an upper bound for all $k$-th eigenvalues without loss of generality, we consider $\lambda_{\mathcal{M},k}>3(1+4c_{\mathcal{M}}).$
@@ 6. Approximation of eigenfunctions
    Let $u^{\prime}=u-\mathbb{P}_{\lambda+a}u.$ Since $Q(u^{\prime})\geq(\lambda+a)\left\lVert u^{\prime}\right\rVert^{2},$
@@
The proof of the above theorem is analogous to that of Theorem 4 in [8], utilizing the results of Lemma 6.4 and Lemma 6.2. ∎